# On the dimension distortion under fractionally smooth mappings

### v1

@@ 4. Intermediate dimensions under fractionally smooth Sobolev mappings
We begin by establishing such an embedding theorem for Hajłasz–Sobolev spaces $\dot{M}^{s,p}$. To facilitate the formulation of the result, we introduce the following piece of notation: Given $Q,b\in(0,\infty)$, $\sigma\in[1,\infty)$, and a ball $B_{0}\subset X$ of radius $R_{0}\in(0,\infty)$, the measure $\mu$ is said to satisfy the $V(\sigma B_{0},Q,b)$ condition [^1] This condition is a slight variation of the one in [39, p. 197]., provided that, for any $x\in X$ and $r\in(0,\sigma R_{0}]$ satisfying $B(x,r)\subset\sigma B_{0}$,
@@
Let $(X,d_{X},\mu)$ be a metric measure space and $(Y,d_{Y})$ be a metric space. Let $\sigma\in(1,\infty)$, $B_{0}\subset X$ be a ball of radius $R_{0}\in(0,\infty)$, and suppose $s,p\in(0,\infty)$ satisfy $p>Q/s$. Assume that the measure $\mu$ satisfies the $V(\sigma B_{0},Q,b)$ condition for some $Q,b\in(0,\infty)$, and suppose $u\in\dot{M}^{s,p}(\sigma B_{0}:Y)$. Then, there exists a set $N\subset X$ with $\mu(N)=0$, such that, for all $x,y\in B_{0}\setminus N$,
(4.2) $d_{Y}(u(x),u(y))\leq Cb^{-1/p}[d_{X}(x,y)]^{s-Q/p}\|u\|_{\dot{M}^{s,p}(\sigma B_{0}:Y)}$
where $C$ is a positive constant depending only on $d_{X}$, $s$, $p$, $Q$, and $\sigma$. In particular, if $(Y,d_{Y})$ is complete then $u$ has a Hölder continuous representative of order sQ/ps-Q/p on $B_{0}$, denoted by $u$, satisfying (4.2) for all $x,y\in B_{0}$.
@@
Choose $g\in\mathcal{D}^{s}(u)\cap L^{p}(\sigma B_{0})$ such that $\|g\|_{L^{p}(\sigma B_{0})}\approx\|u\|_{\dot{M}^{s,p}(\sigma B_{0}:Y)}$. Without loss of generality, we may assume $\int_{\sigma B_{0}}g^{p}\,d\mu>0$. Indeed, if this integral equals zero, then $g=0$ $\mu$-almost everywhere in $\sigma B_{0}$ which, in turn, implies that $u$ is a constant function $\mu$-almost everywhere in $B_{0}$ and the result is obvious in this scenario.
@@
(4.4) $d_{Y}(u(x),u(\xi_{0}))\leq Cb^{-1/p}R_{0}^{s-Q/p}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{1/p},$
@@
It follows from the pointwise inequality (2.12) that $u$ restricted to $E_{k}$ is $2^{k+1}$-Hölder continuous of order $s$, that is,
(4.6) $d_{Y}(u(x),u(y))\leq 2^{k+1}[d_{X}(x,y)]^{s},\ \forall\,x,y\in E_{k}.$
@@
(4.10) $a_{k}:=\left\{\begin{array}[]{ll}\sup\limits_{E_{k}\cap B_{0}}d_{Y}(u,\gamma)\quad&\text{if }\,E_{k}\cap B_{0}\neq\emptyset,\\[15.0pt] \qquad 0&\text{otherwise.}\end{array}\right$.
@@
(4.14) $a_{k}\lesssim b^{-s/Q}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{s/Q}\,\sum_{j=k_{0}}^{k-1}2^{j(1-sp/Q)}+\sup_{E_{k_{0}}}d_{Y}(u,\gamma).$
To see this, let $k\in\mathbb{Z}$ satisfy $k>k_{0}$. Observe that, if $E_{k}\cap B_{0}=\emptyset$, then $a_{k}=0$ due to its definition and (4.14) is trivially true. So we only need to consider the case $E_{k}\cap B_{0}\neq\emptyset$ below. In this case, $a_{k}=\sup_{E_{k}\cap B_{0}}d_{Y}(u,\gamma)$. Now, for any $i\in\{0,1,\ldots,k-k_{0}-1\}$, define
@@
Also, recall that, by (4.6), $u$ restricted to $E_{k-i}$ is $2^{k-i+1}$-Hölder continuous of order $s$. From these and the fact that $x_{k-i},x_{k-(i+1)}\in E_{k-i}$ and $x_{k_{0}}\in E_{k_{0}}$, it follows that
$d_{Y}(u(x_{k}),\gamma)$ $\leq\sum_{i=0}^{k-k_{0}-1}d_{Y}(u(x_{k-i}),u(x_{k-(i+1)}))+d_{Y}(u(x_{k_{0}}),\gamma)$
$\leq\sum_{i=0}^{k-k_{0}-1}2^{k-i+1}[d_{X}(x_{k-i},x_{k-(i+1)})]^{s}+\sup_{E_{k_{0}}}d_{Y}(u,\gamma)$
$\lesssim 2^{s/Q}b^{-s/Q}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{s/Q}\,\sum_{i=0}^{k-k_{0}-1}2^{[k-(i+1)](1-sp/Q)}+\sup_{E_{k_{0}}}d_{Y}(u,\gamma)$
(4.16) $\lesssim b^{-s/Q}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{s/Q}\,\sum_{j=k_{0}}^{k-1}2^{j(1-sp/Q)}+\sup_{E_{k_{0}}}d_{Y}(u,\gamma).$
@@
On the other hand, by Lemma 4.3 we have $\mu(E_{k_{0}})>0$, and so we can choose a point $\xi_{0}\in E_{k_{0}}\subset\sigma B_{0}\setminus N$. Taking $\gamma=u(\xi_{0})$ and applying the Hölder continuity (4.6), gives
$\sup_{E_{k_{0}}}d_{Y}(u,u(\xi_{0}))$ $\leq 2^{k_{0}+1}|\sigma B_{0}|^{s}\leq 2^{k_{0}+1}\left[2\sigma R_{0}\right]^{s}$
@@
With $\gamma=u(\xi_{0})$ as in (4.1) and $\{a_{k}\}_{k\in\mathbb{Z}}$ as in (4.10), observe that since $2^{1-sp/Q}<1$, by (4.14), we find that, for any integer $k>k_{0}$,
(4.19) $a_{k}\lesssim b^{-s/Q}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{s/Q}2^{k_{0}(1-sp/Q)}+\sup_{E_{k_{0}}}d_{Y}(u,u(\xi_{0})).$
@@
(4.20) $a_{k}:=\sup_{E_{k}\cap B_{0}}d_{Y}(u,u(\xi_{0}))\lesssim b^{-1/p}R_{0}^{s-Q/p}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{1/p}.$
Noticing that the right-hand side of (4.20) is a positive constant independent of $k$, by (4.5) and the definition of $a_{k}$ in (4.10), we conclude that the function $d_{Y}\left(u,u(\xi_{0})\right)$ is bounded on $B_{0}\setminus N$ by the right-hand side of (4.20) (modulo a positive constant). This proves (4.4).
@@
These estimates imply that (4.2) holds for every $x,y\in B_{0}\setminus N$. Since $B_{0}\setminus N$ is dense in $B_{0}$ [here we are using the fact that $\mu$ is positive and finite on balls], we can rely on the completeness of $(Y,d_{Y})$ to extend $u$ to a Hölder continuous function of order sQ/ps-Q/p on $B_{0}$, that satisfies (4.2) for every $x,y\in B_{0}$. This completes the proof of Theorem 4.1. ∎
@@
From the proof of Theorem 4.1, if $u\in\dot{M}^{s,p}(X:Y)$ then the null set $N$, and hence also the Hölder continuous representative of $u$, can be chosen independent of the ball $B_{0}$.
@@
    for any $\varepsilon\in(0,s)$, there exists a constant $C\in(0,\infty)$ such that, if $B\subset X$ is a ball with radius $r\in(0,\infty)$, then $\dot{N}^{s}_{p,q}(B:Y)\hookrightarrow\dot{M}^{\varepsilon,p}(B:Y)$, where $\|u\|_{\dot{M}^{\varepsilon,p}(B:Y)}\leq Cr^{s-\varepsilon}\|u\|_{\dot{N}^{s}_{p,q}(B:Y)}$ for all $u\in\dot{N}^{s}_{p,q}(B:Y)$.
@@
Let $(X,d,\mu)$ be a metric measure space, where $\mu$ is locally $Q$-homogeneous for some $Q\in(0,\infty)$, and let $(Y,d_{Y})$ be a metric space. Let $s\in(0,\infty)$, $p\in(Q/s,\infty)$, and $q\in(0,\infty]$, and assume $f:X\to Y$ is a continuous function such that $\|f\|_{\dot{M}^{s}_{p,q}(B:Y)}$ is finite for each fixed ball $B\subset X$. Then, for any compact set $K\subset X$, there exist constants $C_{K}\in[1,\infty)$ and $R_{K}\in(0,\infty)$, both of which are independent of $f$, such that, for all balls $B_{0}:=B(x_{0},R_{0})$ with $x_{0}\in K$ and $R_{0}\in(0,R_{K})$, and one has
@@
Moving on, let $R_{K}:=\tilde{R}_{\text{hom}}(K)/4$ and $B_{0}:=B(x_{0},R_{0})$ a ball with $x_{0}\in K$ and $R_{0}\in(0,R_{K})$. Then (4.22) applied with $B(x,r_{2}):=B(x_{0},2R_{0})=2B_{0}$ implies that the measure $\mu$ satisfies the $V(2B_{0},Q,b)$ condition with $b:=\kappa\mu(2B_{0})(2R_{0})^{-Q}$. On the other hand, since $\|f\|_{\dot{M}^{s}_{p,q}(2B_{0}:Y)}$ is finite, it follows from (1) and (2) in Lemma 4.5 that $\|f\|_{\dot{M}^{s,p}(2B_{0}:Y)}$ is also finite. Combining these observations with Theorem 4.1, we conclude that there exists a set $N\subset X$ with $\mu(N)=0$, such that the inequality in (4.2) holds for all $x,y\in B_{0}\setminus N$. However, since $f$ is continuous and $B_{0}\setminus N$ is dense in $B_{0}$ [here we are using the fact that $\mu$ is positive and finite on balls], we actually have that (4.2) holds for all $x,y\in B_{0}$. This completes the proof of (4.21) for $\dot{M}^{s}_{p,q}$ spaces,
@@
where the implicit constant depends only on $d$, $s$, $\varepsilon$, $p$, and $Q$. With (4.1) in hand, an argument analogous to the one used at the end Theorem 4.1 (keeping in mind that $b:=\kappa\mu(2B_{0})(2R_{0})^{-Q}$) gives that the inequality in (4.21) holds pointwise almost everywhere in $B_{0}$. Now, proceeding as in the case of $\dot{M}^{s}_{p,q}$ spaces, we conclude that (4.21) in fact holds pointwise everywhere in $B_{0}$. This finishes the proof of Corollary 4.6. ∎
@@
for all $R\in(0,\tilde{R}_{\text{hom}}(K)/10)$ and $f\in L^{p}(X)$, where $K^{\prime}\subseteq X$ is any measurable set that contains every ball centered in $K$ with radius at most $\tilde{R}_{\text{hom}}/10$. Indeed, this follows from arguing as in the proof of the Maximal Function Theorem, see, for instance, [44, Chapter 2] or [47, Theorem 3.5.6].
@@
Lemma A.
Let $(X,d_{X},\mu)$ be a metric measure space, where $\mu$ is locally doubling, and fix $t\in[1,p)$ and $\tau\in(0,1)$. For each compact set $K\subset X$ there is a constant $C_{K}^{\prime}\geq 1$ and a radius $R_{K}^{\prime}>0$ such that for all nonnegative functions $g\in L^{p}(X)$,
@@
Lemma A is a corollary of the Maximal Function Theorem, see [44, Chapter 2] and [12, Lemma 3.3] for a proof.
Suppose $f:X\rightarrow Y$ is continuous with a finite Hajłasz–Triebel–Lizorkin semi-norm $\|f\|_{\dot{M}^{s}_{p,q}(X:Y)}$, or finite a Hajłasz–Besov semi-norm $\|f\|_{\dot{N}^{s}_{p,q}(X:Y)}$. Note that for every $t\in[1,p)$ and every ball $B\subset X$ we have $\|f\|_{\dot{M}^{s}_{t,q}(B:Y)}$ or $\|f\|_{\dot{N}^{s}_{t,q}(B:Y)}$ is finite. We plan on using Corollary 4.6 and Lemma A to show that $f$ belongs to the appropriate compactly Hölder class, which is enough to achieve (1.4) and (1.5).
@@
Let $E\subset X$ be compact and $\varepsilon\in(0,1)$. Note that the centers of balls that cover $E$ in the definition of compactly Hölder mappings do not necessarily lie in $E$, while it is a requirement for the inequalities in Corollary 4.6 and Lemma A. Thus, we need to apply these properties to a potentially larger compact set than $E$. Take any point $x_{E}\in E$. Then a straightforward calculation shows that
@@
Thus, if $K$ is the closure of $B(x_{E},|E|+1/2)$, and $\mathcal{B}$ is a covering of $E$ by balls of radius at most $1/10$, then all elements of $\mathcal{B}$ lie entirely in $K$, along with their centers. Our plan is to apply Corollary 4.6 and Lemma A with the set $K$, which is compact by virtue of $X$ being proper. With this goal in mind, consider the radius
@@
where $R_{K},R_{K}^{\prime}$ are as in Corollary 4.6 and Lemma A, respectively, and $R_{\text{hom}}(K):=\tilde{R}_{\text{hom}}(K)/3$ with $\tilde{R}_{\text{hom}}(K)$ as in the local $Q$-homogeneity condition (2.14) for the compact set $K$.
Suppose $\{B(x_{i},r)\}_{i\in I}$ is a cover of $E$ with $r<r_{E}$ and $B(x_{i},\varepsilon r)\cap B(x_{j},\varepsilon r)=\emptyset$ for all distinct $i,j\in I$. We will show that an inequality of the form (2.9) holds. Let us first consider the case when $\|f\|_{\dot{M}^{s}_{p,q}(X:Y)}$ is finite. By (1) and (2) in Lemma 4.5, we can assume $\|f\|_{\dot{M}^{s,p}(X:Y)}$ is finite. Let $g\in\mathcal{D}^{s}(f)\cap L^{p}(X)$ and fix $t\in(\min\{1,Q/s\},p)$. Then it follows from Hölder’s inequality that $\|f\|_{\dot{M}^{t,p}(B:Y)}$ is finite for each fixed ball $B\subset X$, where the pointwise restriction of $g$ to $B$ serves as an $s$-gradient of $f$. By this, (4.28), and Corollary 4.6, we have that there exists a constant $C_{k}\in[1,\infty)$, which is independent of $f$, such that, for all balls $B_{i}:=B(x_{i},r)$, the following inequality holds
@@
where $\tilde{g}:=M_{r_{E}}(g^{t})\in L^{p/t}(K)\subset L^{1}(K)$ by (4.25). Since the measure $\mu$ is locally $Q$-homogeneous, by (2.15) and (4.28), the quotient $\frac{|B_{i}|^{Q}}{\mu(\varepsilon B_{i})}$ is bounded from above by $\tilde{C}:=C_{\text{hom}}(K)\left(\frac{1}{\varepsilon}\right)^{Q}$. Hence, due to $B(x_{i},\varepsilon r)\cap B(x_{j},\varepsilon r)=\emptyset$ and

### v2

@@ 4. Intermediate dimensions under fractionally smooth Sobolev mappings
We begin by establishing such an embedding theorem for Hajłasz–Sobolev spaces $\dot{M}^{s,p}$. To facilitate the formulation of the result, we introduce the following piece of notation: Given $Q,b\in(0,\infty)$, $\sigma\in[1,\infty)$, and a ball $B_{0}\subset X$ of radius $R_{0}\in(0,\infty)$, the measure $\mu$ is said to satisfy the $V(\sigma B_{0},Q,b)$ condition [^1] This condition is a slight variation of the one in [41, p. 197]., provided that, for any $x\in X$ and $r\in(0,\sigma R_{0}]$ satisfying $B(x,r)\subset\sigma B_{0}$,
@@
Let $(X,d_{X},\mu)$ be a metric measure space and $(Y,d_{Y})$ be a metric space. Let $\sigma\in(1,\infty)$, $B_{0}\subset X$ be a ball of radius $R_{0}\in(0,\infty)$, and suppose $s,p,Q\in(0,\infty)$ satisfy $sp>Q$. Assume that the measure $\mu$ satisfies the $V(\sigma B_{0},Q,b)$ condition for some $b\in(0,\infty)$, and suppose $f\in\dot{M}^{s,p}(\sigma B_{0}:Y)$. Then, there exists a set $N\subset X$ with $\mu(N)=0$, such that, for all $x,y\in B_{0}\setminus N$,
(4.2) $d_{Y}(f(x),f(y))\leq Cb^{-1/p}[d_{X}(x,y)]^{s-Q/p}\|f\|_{\dot{M}^{s,p}(\sigma B_{0}:Y)},$
where $C\in(0,\infty)$ is a constant depending only on $s$, $p$, $Q$, and $\sigma$. In particular, if $(Y,d_{Y})$ is complete then $f$ has a Hölder continuous representative of order sQ/ps-Q/p on $B_{0}$, denoted by $u$, satisfying (4.2) for all $x,y\in B_{0}$.
@@
Choose $g\in D^{s}(f)\cap L^{p}(\sigma B_{0})$ such that $\|g\|_{L^{p}(\sigma B_{0})}\approx\|f\|_{\dot{M}^{s,p}(\sigma B_{0}:Y)}$. Without loss of generality, we may assume $\int_{\sigma B_{0}}g^{p}\,d\mu>0$. Indeed, if this integral equals zero, then $g=0$ $\mu$-almost everywhere in $\sigma B_{0}$ which, in turn, implies that $u$ is a constant function $\mu$-almost everywhere in $B_{0}$ and the result is obvious in this scenario.
@@
(4.4) $d_{Y}(f(x),f(\xi_{0}))\leq Cb^{-1/p}R_{0}^{s-Q/p}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{1/p},$
@@
It follows from the pointwise inequality (2.11) that $f$ restricted to $E_{k}$ is $2^{k+1}$-Hölder continuous of order $s$, that is,
(4.6) $d_{Y}(f(x),f(y))\leq 2^{k+1}[d_{X}(x,y)]^{s},\ \forall\,x,y\in E_{k}.$
@@
(4.10) $a_{k}:=\left\{\begin{array}[]{ll}\sup\limits_{E_{k}\cap B_{0}}d_{Y}(f,\gamma)\quad&\text{if }\,E_{k}\cap B_{0}\neq\emptyset,\\[15.0pt] \qquad 0&\text{otherwise.}\end{array}\right$.
@@
(4.14) $a_{k}\lesssim b^{-s/Q}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{s/Q}\,\sum_{j=k_{0}}^{k-1}2^{j(1-sp/Q)}+\sup_{E_{k_{0}}}d_{Y}(f,\gamma).$
To see this, let $k\in\mathbb{Z}$ satisfy $k>k_{0}$. Observe that, if $E_{k}\cap B_{0}=\emptyset$, then $a_{k}=0$ due to its definition and (4.14) is trivially true. So we only need to consider the case $E_{k}\cap B_{0}\neq\emptyset$ below. In this case, $a_{k}=\sup_{E_{k}\cap B_{0}}d_{Y}(f,\gamma)$. Now, for any $i\in\{0,1,\ldots,k-k_{0}-1\}$, define
@@
Also, recall that, by (4.6), $f$ restricted to $E_{k-i}$ is $2^{k-i+1}$-Hölder continuous of order $s$. From these and the fact that $x_{k-i},x_{k-(i+1)}\in E_{k-i}$ and $x_{k_{0}}\in E_{k_{0}}$, it follows that
$d_{Y}(f(x_{k}),\gamma)$ $\leq\sum_{i=0}^{k-k_{0}-1}d_{Y}(f(x_{k-i}),f(x_{k-(i+1)}))+d_{Y}(f(x_{k_{0}}),\gamma)$
$\leq\sum_{i=0}^{k-k_{0}-1}2^{k-i+1}[d_{X}(x_{k-i},x_{k-(i+1)})]^{s}+\sup_{E_{k_{0}}}d_{Y}(f,\gamma)$
$\lesssim 2^{s/Q}b^{-s/Q}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{s/Q}\,\sum_{i=0}^{k-k_{0}-1}2^{[k-(i+1)](1-sp/Q)}+\sup_{E_{k_{0}}}d_{Y}(f,\gamma)$
(4.16) $\lesssim b^{-s/Q}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{s/Q}\,\sum_{j=k_{0}}^{k-1}2^{j(1-sp/Q)}+\sup_{E_{k_{0}}}d_{Y}(f,\gamma).$
@@
On the other hand, by Lemma 4.3 we have $\mu(E_{k_{0}})>0$, and so we can choose a point $\xi_{0}\in E_{k_{0}}\subset\sigma B_{0}\setminus N$. Taking $\gamma:=f(\xi_{0})$ and applying the Hölder continuity (4.6), gives
$\sup_{E_{k_{0}}}d_{Y}(f,f(\xi_{0}))$ $\leq 2^{k_{0}+1}|\sigma B_{0}|^{s}\leq 2^{k_{0}+1}\left[2\sigma R_{0}\right]^{s}$
@@
With $\gamma=f(\xi_{0})$ as in (4.1) and $\{a_{k}\}_{k\in\mathbb{Z}}$ as in (4.10), observe that since $2^{1-sp/Q}<1$, by (4.14), we find that, for any integer $k>k_{0}$,
(4.19) $a_{k}\lesssim b^{-s/Q}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{s/Q}2^{k_{0}(1-sp/Q)}+\sup_{E_{k_{0}}}d_{Y}(f,f(\xi_{0})).$
@@
(4.20) $a_{k}:=\sup_{E_{k}\cap B_{0}}d_{Y}(f,f(\xi_{0}))\lesssim b^{-1/p}R_{0}^{s-Q/p}\left(\,\int_{\sigma B_{0}}g^{p}\,d\mu\right)^{1/p}.$
Noticing that the right-hand side of (4.20) is a positive constant independent of $k$, by (4.5) and the definition of $a_{k}$ in (4.10), we conclude that the function $d_{Y}\left(f,f(\xi_{0})\right)$ is bounded on $B_{0}\setminus N$ by the right-hand side of (4.20) (modulo a positive constant). This proves (4.4).
@@
Since $\|g\|_{L^{p}(\sigma B_{0})}\approx\|f\|_{\dot{M}^{s,p}(\sigma B_{0}:Y)}$, these estimates imply that (4.2) holds for every $x,y\in B_{0}\setminus N$. Since $B_{0}\setminus N$ is dense in $B_{0}$ [here we are using the fact that $\mu$ is positive and finite on balls], we can rely on the completeness of $(Y,d_{Y})$ to extend $f$ to a Hölder continuous function of order sQ/ps-Q/p on $B_{0}$, that satisfies (4.2) for every $x,y\in B_{0}$. This completes the proof of Theorem 4.1. ∎
@@
It is easy to see from the proof of Theorem 4.1 that if $f:\sigma B_{0}\to Y$ is a continuous mapping such that $\|f\|_{\dot{M}^{s,p}(\sigma B_{0}:Y)}<\infty$ then (4.2) holds for all $x,y\in B_{0}$, without needing to assume that $(Y,d_{Y})$ is complete.
@@
    for any $\varepsilon\in(0,s)$, there exists a constant $C\in(0,\infty)$ such that, if $B\subset X$ is a ball with radius $r\in(0,\infty)$, then $\dot{N}^{s}_{p,q}(B:Y)\hookrightarrow\dot{M}^{\varepsilon,p}(B:Y)$, where $\|f\|_{\dot{M}^{\varepsilon,p}(B:Y)}\leq Cr^{s-\varepsilon}\|f\|_{\dot{N}^{s}_{p,q}(B:Y)}$ for all $f\in\dot{N}^{s}_{p,q}(B:Y)$.
@@
Let $(X,d,\mu)$ be a metric measure space, where $\mu$ is locally $Q$-homogeneous for some $Q\in(0,\infty)$, and let $(Y,d_{Y})$ be any metric space. Suppose $s\in(0,\infty)$, $p\in(Q/s,\infty)$, and $q\in(0,\infty]$, and assume $f:X\to Y$ is a continuous function. Then, for any compact set $K\subset X$, there exist constants $C_{K}\in[1,\infty)$ and $R_{K}\in(0,\infty)$, both of which are independent of $f$, such that, for all balls $B_{0}:=B(x_{0},R_{0})$ with $x_{0}\in K$ and $R_{0}\in(0,R_{K})$ for which $\|f\|_{\dot{M}^{s}_{p,q}(2B_{0}:Y)}<\infty$, one has
@@
Remark 4.7.
In Corollary 4.6, the specific dilation factor $2$ of the ball $B_{0}$ is not essential and can be replaced by any $\sigma\in(1,\infty)$.
@@
Moving on, let $R_{K}:=\tilde{R}_{\text{hom}}(K)/4$ and $B_{0}:=B(x_{0},R_{0})$ a ball with $x_{0}\in K$ and $R_{0}\in(0,R_{K})$. Then (4.22) applied with $B(x,r_{2}):=B(x_{0},2R_{0})=2B_{0}$ implies that the measure $\mu$ satisfies the $V(2B_{0},Q,b)$ condition with $b:=\kappa\mu(2B_{0})(2R_{0})^{-Q}$. On the other hand, it follows from (1) and (2) in Lemma 4.5 that $\|f\|_{\dot{M}^{s,p}(2B_{0}:Y)}\lesssim\|f\|_{\dot{M}^{s}_{p,q}(2B_{0}:Y)}$ which, combined with the assumption that $\|f\|_{\dot{M}^{s}_{p,q}(2B_{0}:Y)}$ is finite, gives that $\|f\|_{\dot{M}^{s,p}(2B_{0}:Y)}$ is also finite. Given these observations, by Theorem 4.1 and Remark 4.4 we conclude that (4.2) holds for all $x,y\in B_{0}$ with $b:=\kappa\mu(2B_{0})(2R_{0})^{-Q}$. This, together with $\|f\|_{\dot{M}^{s,p}(2B_{0}:Y)}\lesssim\|f\|_{\dot{M}^{s}_{p,q}(2B_{0}:Y)}$, completes the proof of (4.21) for $\dot{M}^{s}_{p,q}$ spaces.
@@
where the implicit constant depends only on $d$, $s$, $\varepsilon$, $p$, and $Q$. With (4.1) in hand, an argument analogous to the one used at the end Theorem 4.1 (keeping in mind that $b:=\kappa\mu(2B_{0})(2R_{0})^{-Q}$) gives that the inequality in (4.21) holds pointwise almost everywhere in $B_{0}$. Given that $f$ is continuous, we conclude that (4.21) in fact holds pointwise everywhere in $B_{0}$. This finishes the proof of Corollary 4.6. ∎
@@
for all $R\in(0,\tilde{R}_{\text{hom}}(K)/10)$ and $f\in L^{p}(K^{\prime})$, where $K^{\prime}\subset X$ is any measurable set that contains every ball centered in $K$ with radius at most $\tilde{R}_{\text{hom}}/10$. Indeed, this follows from arguing as in the proof of the Maximal Function Theorem, see, for instance, [46, Chapter 2] or [49, Theorem 3.5.6].
@@
Lemma 4.8.
Let $(X,d_{X},\mu)$ be a proper metric measure space, where $\mu$ is locally doubling, and fix $p\in(0,\infty)$, $t\in(0,p)$, and $\tau\in(0,1)$. For each compact set $K\subset X$ there is a constant $C_{K}^{\prime}\in[1,\infty)$ and a radius $R_{K}^{\prime}\in(0,\infty)$ such that for all nonnegative functions $g\in L^{p}_{{\scriptstyle{loc}}}(X)$,
@@
Lemma 4.8 is a corollary of the Maximal Function Theorem, see [46, Chapter 2] and [12, Lemma 3.3] for a proof.
We plan on using Corollary 4.6 and Lemma 4.8 to show that $f$ belongs to the appropriate compactly Hölder class, which is enough to achieve (1.4) and (1.5).
@@
Let $E\subset X$ be compact and $\varepsilon\in(0,1)$. Note that the centers of balls that cover $E$ in the definition of compactly Hölder mappings do not necessarily lie in $E$, while it is a requirement for the inequalities in Corollary 4.6 and Lemma 4.8. Thus, we need to apply these properties to a potentially larger compact set than $E$. Take any point $x_{E}\in E$. Then a straightforward calculation shows that
@@
Thus, if $K$ is the closure of $B(x_{E},|E|+1/2)$, and $\mathcal{B}$ is a covering of $E$ by balls of radius at most $1/10$, then all elements of $\mathcal{B}$ lie entirely in $K$, along with their centers.
Our plan is to apply Corollary 4.6 and Lemma 4.8 with the set $K$, which is compact by virtue of $X$ being proper. With this goal in mind, consider the radius
@@
where $R_{K},R_{K}^{\prime}$ are as in Corollary 4.6 and Lemma 4.8, respectively, and $R_{\text{hom}}(K):=\tilde{R}_{\text{hom}}(K)/3$ with $\tilde{R}_{\text{hom}}(K)$ as in the local $Q$-homogeneity condition (2.13) for the compact set $K$. With a view of applying (4.25) later in the proof, let $K^{\prime}\subset X$ be any measurable set that contains every ball centered in $K$ with radius at most $2r_{E}$. Since $X$ is proper, we can choose $K^{\prime}$ to be compact.
Suppose $\{B(x_{i},r_{i})\}_{i\in I}$ is a cover of $E$ with $r_{i}<r_{E}$ and $B(x_{i},\varepsilon r_{i})\cap B(x_{j},\varepsilon r_{j})=\emptyset$ for all distinct $i,j\in I$. We will show that an inequality of the form (2.9) holds. Let us first consider the case when $\|f\|_{\dot{M}^{s}_{p,q}(X:Y)}$ is finite. By (1) and (2) in Lemma 4.5, we can assume $\|f\|_{\dot{M}^{s,p}(X:Y)}$ is finite. Then $\|f\|_{\dot{M}^{s,p}(K^{\prime}:Y)}$ is finite and so there exists $g\in D^{s}(f)\cap L^{p}(K^{\prime})$. Fix $t\in(Q/s,p)$. By Hölder’s inequality and the fact that each $2B_{i}:=B(x_{i},2r_{i})\subset K^{\prime}$, we have $\|f\|_{\dot{M}^{t,p}(2B_{i}:Y)}$ is finite for each fixed $B_{i}$, where the pointwise restriction of $g:K^{\prime}\to[0,\infty]$ to $2B_{i}$ serves as an $s$-gradient of $f$. By this, (4.28), and Corollary 4.6, we have that there exists a constant $C_{K}\in[1,\infty)$, which is independent of $f$, such that, for all balls $B_{i}:=B(x_{i},r_{i})$, the following inequality holds
@@
where $\tilde{g}:=M_{r_{E}}(g^{t})\in L^{p/t}(K)\subset L^{1}(K)$ by (4.25). Here we view $g$ as a locally $p$-integrable function defined on $X$ by extending it by zero outside its original domain. Since the measure $\mu$ is locally $Q$-homogeneous, by (2.14) and (4.28), the quotient $\frac{|B_{i}|^{Q}}{\mu(\varepsilon B_{i})}$ is bounded from above by $\tilde{C}:=C_{\text{hom}}(K)\left(\frac{1}{\varepsilon}\right)^{Q}$. Hence, due to $B(x_{i},\varepsilon r_{i})\cap B(x_{j},\varepsilon r_{j})=\emptyset$ and